Accounting

Aircard Corporation tracks the number of units purchased and sold throughout each accounting period but applies its inventory costing method at the end of each period as if it uses a periodic inventory system. The following are the transactions for the month of July.

 

  UnitsUnit Cost
July 1Beginning Inventory2,000$ 25
July 5Sold1,000
July 13Purchased6,00029
July 17Sold3,000
July 25Purchased8,00031
July 27Sold5,000

 

Calculate the cost of goods available for sale, ending inventory, and cost of goods sold if Aircard uses (a) FIFO, (b) LIFO, or (c) weighted average cost. (Round "Cost per Unit" to 2 decimal places.)
Answer #1

Units in Beginning Inventory = 2,000

Number of units purchased = 6,000 + 8,000
Number of units purchased = 14,000

Number of units sold = 1,000 + 3,000 + 5,000
Number of units sold = 9,000

Number of units Available for Sale = Units in Beginning Inventory + Number of units purchased
Number of units Available for Sale = 2,000 + 14,000
Number of units Available for Sale = 16,000

Units in Ending Inventory = Number of units Available for Sale – Number of units Sold
Units in Ending Inventory = 16,000 – 9,000
Units in Ending Inventory = 7,000

Cost of Goods Available for Sale = 2,000 * $25 + 6,000 * $29 + 8,000 * $31

Cost of Goods Available for Sale = $472,000

LIFO:

Cost of Goods Sold = 8,000 * $31 + 1,000 * $29
Cost of Goods Sold = $277,000

Ending Inventory = 5,000 * $29 + 2,000 * $25
Ending Inventory = $195,000

FIFO:

Cost of Goods Sold = 2,000 * $25 + 6,000 * $29 + 1,000 * $31
Cost of Goods Sold = $255,000

Ending Inventory = 7,000 * $31
Ending Inventory = $217,000

Weighted-average Cost:

Cost of Goods Available for Sale = $744,000
Number of units Available for Sale = 16,000

Weighted Average Cost per unit = $744,000 / 16,000
Weighted Average Cost per unit = $46.50

Cost of Goods Sold = 9,000 * $46.50
Cost of Goods Sold = $418,500

Ending Inventory = 7,000 * $46.50
Ending Inventory = $325,500


answered by: anonymous